# Learning an Ensemble Token from Task-driven Priors in Facial Analysis

### v2

@@ 1 Introduction
Facial analysis exhibits task-specific feature variations; conversely, a unified feature representation derived from diverse features yields robust performance across facial analysis tasks. Deep learning neural networks represent spatial and semantic features from a single facial image. The Convolutional Neural Networks (CNNs) enable spatial information representation, and more recently, the Vision Transformers (ViTs) facilitate the semantic information representation at the patch level. Notably, the transformer’s attention mechanism enables robust representation even when inputs are scattered across disparate spaces or modalities [13, 40].
@@
The emergence of general models such as DINOv2 [25] and SAM [14] capable of handling tasks has significantly advanced interpretability within the computer vision. These general models have demonstrated significant achievements in image classification, segmentation, and detection, and as well as handling multi-tasks within a single encoder. Furthermore, linear probing of pre-trained models with unified feature representation has shown promising results [10, 16, 27, 12]. In the areas of multi-modality, CLIP [29, 32] and attention mechanisms [38, 5] have demonstrated promising results, confirming that unified feature representation significantly enhances model’s interpretability.
As noted, the capacity of single and multi-task learning in conventional methodologies has stepped up the visual interpretability. However, there remains a paucity of research that preserves the unified feature representation or encoder’s priors on single task learning during the training phase. Although task-oriented fine-tuning can be achieved through ensemble learning [8, 21, 22], there is still a challenge in that the computational cost escalates proportionally to both the number of tasks and the encoder model’s capacity. In this research, we introduce a methodology for learning knowledge tokens by leveraging the attention mechanisms based on the task priors of pre-trained models in facial analysis tasks. The principal contributions of this study are shown in  1 and as follows:
@@ 2.1 Leveraging Pre-trained Priors
Ensemble learning. The capacity to train specialized models for individual tasks has enabled ensemble methods to achieve robust performance through the exploitation of task-specific priors. These methods range from classical techniques like bagging, boosting, and stacking, to more sophisticated fusion approaches such as majority voting and deep learning-based methodologies [28, 2, 17, 20]. Nevertheless, a fundamental limitation of ensemble methodologies resides in their computational cost, which is directly proportional to both model capacity and ensemble cardinality. This computational cost can preclude their applicability in real-time facial analysis scenarios as shown in  2. Although recent studies [42, 39] have explored strategies to mitigate the computational cost of model cardinality, the challenge of developing training frameworks that are simultaneously efficient, learnable, and adaptable across high feature spaces remains unresolved.
Low rank adaptation (LoRA). The LoRA [11] was introduced to enable parameter-efficient fine-tuning (PEFT), primarily to reduce the computational cost and hardware demands of training parameter-heavy models. Within the computer vision, LoRA has also proven valuable as a PEFT strategy, especially for diffusion models [9, 7]. Following LoRA, ControlNet-based [44] adaptation of foundation models has demonstrated significant efficacy in domain and task adaptation even on multi-modality face analysis tasks [15, 36]. While ControlNet has established its capacity to preserve pre-trained priors within task-specific models, the potential for ensembling encoders to achieve controllable training warrants further investigation.
@@ 2.2 Feature Tokenization on Facial Analysis
The attention mechanism [5] and CLIP [26] have become a cornerstones within the computer vision, effectively capturing both the semantic tokens and the spatial hierarchies within single and multi-modal data. In facial analysis, the HOTformer [34] introduced a novel approach employing lightweight CNNs guided by holistic and atomic tokens. This hybrid token of HOTformer effectively integrates facial semantics and outperforms the accuracy on face recognition. The preliminary work [45] proposed TokenFace, a transformer-based 3D face reconstruction model with a learnable facial component tokens encompassing shape, expressions, jaw pose, camera pose, texture, and lighting. This multi-modality approach effectively integrates the interdependencies between facial characteristics and image token embeddings derived from ViT within the single task. The POSTER++ [19] advanced the state-of-the-art in facial emotion recognition by replacing traditional pyramid architectures with a more efficient design, achieving performance with minimal computational cost. This method leverages cross-attention to fuse landmark priors and image features. The FaceXFormer [24] fuses the hierarchical representation extracted from CNN blocks. Its task-specific tokenization scheme facilitates multi-task learning within a single encoder architecture, effectively utilizing both self-attention and cross-attention mechanisms.
@@ 3 Methodology
Inspired by the object token in DETR [1] and mask token in MaskFormer [3, 4], we propose learnable knowledge tokens corresponding to learnable query features in preliminaries. DETR introduced an anchor-free semantic representation leveraging object tokens. Subsequently, Mask2Former utilized masked attention to incorporate both semantic and spatial information into learnable query features. Motivated by the concept of learnable query embeddings, we introduce a knowledge token designed to capture diverse embeddings derived from the priors of a pre-trained model. While conceptually similar to a module that fuses information acquired from an ensemble, this knowledge token is specifically tailored to the canonical task, selectively attending to and learning pertinent semantic information.
@@ 3.1 Acquiring Prior Tokens
In the preliminary stage, a set of N pre-trained encoders $E=\{\varepsilon_{1},\varepsilon_{2},...,\varepsilon_{n-1},\varepsilon_{c}\}$ corresponding to the expertise of each task is prepared. These encoders could be employed either CNN or ViT architectures, or even a hybrid thereof dictated by task-specific priors. Subsequent to the preparation of $E$, a canonical task branch is selected for task-specific training. As illustrated in  3, face landmark detection serves as the canonical task in the presented example. During training stage, supervision learning is exclusively derived from the ground truth $Y=\{y_{1},y_{2},...,y_{n-1},y_{c}\}$ corresponding to the designated canonical task; ground truth of the N-1 remaining tasks $Y\setminus\{y_{c}\}$ is explicitly excluded from the training process.
@@
Figure 4: Detailed structure of KT-Adapter. The KT-Adapter is trained on a restricted information $T_{\text{mask}}$ to enrich the $e_{c}$ in  3. During the training phase, the MHSA layer processes $T_{\text{mask}}$, which has been subjected to subsampling via the drop mask. During inference, the MHSA layer only processes $\{t_{k},t_{c}\}$, and the canonical mask is omitted.
@@ 3.2 KT-Adapter
We introduce the KT-Adapter architecture to enable the learning of knowledge tokens. KT-Adapter functions to adapt a knowledge token, drawing inspiration from class token within attention mechanisms and object tokens in DETR [1], to a designated canonical task through the utilization of prior tokens. Leveraging a self-attention mechanism, KT-Adapter facilitates the learning of both knowledge token $t_{k}$ in KT-Adapter and canonical prior $t_{c}$ derived from canonical encoder $e_{c}$. By focusing on the canonical task, it effectively reduces redundancy among the remaining prior tokens, which represent task-optimized embeddings from pre-trained tasks, thereby generating representations $\hat{t}_{k}$ and $\hat{t}_{c}$. Distinct from ensemble and ControlNet architectures, the KT-Adapter architecture eliminates the requirement for supplementary priors during inference, thereby enabling computationally efficient inference based solely on the learned knowledge token.
@@ 3.3 Drop mask and Canonical mask
In  4, we initialize the $t_{k}$, a learnable query feature, and subsequently concatenate it with the $T$. Drawing inspiration from MaskFormer [3] and dropout [33], random subsampling within the feature representation not only mitigates overfitting but also rectifies global representations and potential misalignment in the token’s forward pass between training and inference phases, specifically within KT-Adapter. We introduce a stochastic drop mask $m$ denoted as:
@@
We employed multi-head self-attention (MHSA) for learning knowledge tokens. The subsampled token set $T_{{masked}}$ in  2 is presented as input to the conventional attention mechanism:
@@ 3.4 Decoder and Inference
The $\hat{t}_{k}$ and $\hat{t}_{c}$ derived from the KT-Adapter are subjected to element-wise addition with the spatial embedding information $e_{i}$. The feature embedding $\hat{e}_{c}$ is subsequently provided as input to the canonical decoder $d_{c}$. While a set of N decoders $D=\{d_{1},d_{2},...,d_{n-1},d_{c}\}$ is potentially learnable for other tasks, only the task-specific decoder $d_{c}$ undergoes fine-tuning to accommodate $\hat{e}_{c}$. This canonical, or selective, fine-tuning strategy restricts the trainable parameters to the KT-Adapter and the canonical decoder, thereby significantly reducing the computational resources required for both training and inference. Each decoder comprises a light multi-layer perceptron architecture tailored for classification and regression tasks. For the loss function, WingLoss [6] was employed for regression tasks, including age estimation and landmark detection, and cross-entropy loss was employed for classification tasks, including face emotion recognition and face recognition.
@@
During inference, only $t_{c}$ derived from $e_{c}$ are presented as input to the KT-Adapter as shown in  5. While the information density is lower than training, $t_{k}$ retains a sophisticated and entangled representation of the priors acquired during the training phase. The KT-Adapter’s training on ”thinned” information, facilitated by the drop mask, ensures that the inference process exhibits a high degree of fidelity to the training phase. To facilitate the application of positional embeddings to $T_{\text{masked}}$, the task index corresponding to each encoder is preserved throughout the training phase. This preservation guarantees the consistent forward pass instead of positional embeddings to corresponding token’s locations, even in scenarios where elements from $T\setminus{t_{c}}$ are absent in inference phase. Canonical masking operations are omitted during inference as it identically equals to $\{\hat{t}_{k},\hat{t}_{c}\}$. In consequence, the exclusion of other $e_{i}$ results in a substantial reduction in computational overhead compared to knowledge learning, with the computational savings scaling linearly with the number of encoders.
@@ 4.1 Implementation Details
The experimental evaluations were conducted on 2 NVIDIA Tesla H100 GPUs. For the backbone network architecture, MobileOne [37] was selected to assess real-time performance on resource-constrained platforms, such as embedded systems and handheld devices, while ConvNeXt-b (CNN-based) [41] and Swin-b (ViT-based) [18] were employed to evaluate performance on desktop-level hardware. All input images were resized to $\times224$ pixels, with facial regions cropped using a YOLO-based [30] face detector. No facial alignment procedures were applied; however, task-specific pre-processing pipelines were implemented to ensure robust generalization by tasks. Specifically, for age estimation and face recognition tasks, the facial region was cropped to encompass both the facial area and hair, whereas for face landmark detection and face emotion recognition, a more tightly cropped region was adapted. A uniform batch size of 32 was maintained across all tasks. Given the head-tuning paradigm employed, which is analogous to linear probing, a learning rate of 1e-5 was adopted in trainable layers with warm-up cosine scheduler. In terms of computational cost, the KT-Adapter exhibited 21.02 MFLOPs during the inference phase, with a token dimension size of 1024. During the training phase, the computational cost increases proportionally with the number of tasks.

### v3

@@ 1 Introduction
Facial analysis exhibits task-specific feature variations; conversely, a unified feature representation derived from diverse features yields robust performance across facial analysis tasks. Deep learning neural networks represent spatial and semantic features from a single facial image. The Convolutional Neural Networks (CNNs) enable spatial information representation, and more recently, the Vision Transformers (ViTs) facilitate the semantic information representation at the patch level. Notably, the transformer’s attention mechanism enables robust representation even when inputs are scattered across disparate spaces or modalities [Kirillov_2023_ICCV, Wei_2020_CVPR].
@@
The emergence of general models such as DINOv2 [oquab2024dinov2learningrobustvisual] and SAM [kirillov2023segment] capable of handling tasks has significantly advanced interpretability within the computer vision. These general models have demonstrated significant achievements in image classification, segmentation, and detection, and as well as handling multi-tasks within a single encoder. Furthermore, linear probing of pre-trained models with unified feature representation has shown promising results [he2021maskedautoencodersscalablevision, Liang_2022_CVPR, Pham_2024_WACV, Huang_2024_CVPR]. In the areas of multi-modality, CLIP [radford2021learningtransferablevisualmodels, Shen_2023_ICCV] and attention mechanisms [vaswani2023attentionneed, dosovitskiy2021imageworth16x16words] have demonstrated promising results, confirming that unified feature representation significantly enhances model’s interpretability.
As noted, the capacity of single and multi-task learning in conventional methodologies has stepped up the visual interpretability. However, there remains a paucity of research that preserves the unified feature representation or encoder’s priors on single task learning during the training phase. Although task-oriented fine-tuning can be achieved through ensemble learning [Ganaie_2022, 9893798, MOHAMMED2023757], there is still a challenge in that the computational cost escalates proportionally to both the number of tasks and the encoder model’s capacity. In this research, we introduce a methodology for learning knowledge tokens by leveraging the attention mechanisms based on the task priors of pre-trained models in facial analysis tasks. The principal contributions of this study are shown in Fig. 1 and as follows:
@@ 2.1 Leveraging Pre-trained Priors
Ensemble learning. The capacity to train specialized models for individual tasks has enabled ensemble methods to achieve robust performance through the exploitation of task-specific priors. These methods range from classical techniques like bagging, boosting, and stacking, to more sophisticated fusion approaches such as majority voting and deep learning-based methodologies [Qian_2023_ICCV, Chen_2023_ICCV, 10.1007/978-3-031-73021-4_7, 10.1007/978-3-031-73209-6_18]. Nevertheless, a fundamental limitation of ensemble methodologies resides in their computational cost, which is directly proportional to both model capacity and ensemble cardinality. This computational cost can preclude their applicability in real-time facial analysis scenarios as shown in Fig. 2. Although recent studies [pmlr-v162-wortsman22a, 10444954] have explored strategies to mitigate the computational cost of model cardinality, the challenge of developing training frameworks that are simultaneously efficient, learnable, and adaptable across high feature spaces remains unresolved.
Low rank adaptation (LoRA). The LoRA [hu2021loralowrankadaptationlarge] was introduced to enable parameter-efficient fine-tuning (PEFT), primarily to reduce the computational cost and hardware demands of training parameter-heavy models. Within the computer vision, LoRA has also proven valuable as a PEFT strategy, especially for diffusion models [10.1007/978-3-031-73661-2_10, 10.1007/978-3-031-72684-2_11]. Following LoRA, ControlNet-based [controlnet] adaptation of foundation models has demonstrated significant efficacy in domain and task adaptation even on multi-modality face analysis tasks [10.1007/978-3-031-72649-1_10, varanka2024localizedfinegrainedcontrolfacial]. While ControlNet has established its capacity to preserve pre-trained priors within task-specific models, the potential for ensembling encoders to achieve controllable training warrants further investigation.
@@ 2.2 Feature Tokenization on Facial Analysis
The attention mechanism [dosovitskiy2021imageworth16x16words] and CLIP [Ou_2023_ICCV] have become a cornerstones within the computer vision, effectively capturing both the semantic tokens and the spatial hierarchies within single and multi-modal data. In facial analysis, the HOTformer [SU2023109443] introduced a novel approach employing lightweight CNNs guided by holistic and atomic tokens. This hybrid token of HOTformer effectively integrates facial semantics and outperforms the accuracy on face recognition. The preliminary work [Zhang_2023_ICCV] proposed TokenFace, a transformer-based 3D face reconstruction model with a learnable facial component tokens encompassing shape, expressions, jaw pose, camera pose, texture, and lighting. This multi-modality approach effectively integrates the interdependencies between facial characteristics and image token embeddings derived from ViT within the single task. The POSTER++ [MAO2025110951] advanced the state-of-the-art in facial emotion recognition by replacing traditional pyramid architectures with a more efficient design, achieving performance with minimal computational cost. This method leverages cross-attention to fuse landmark priors and image features. The FaceXFormer [narayan2024facexformerunifiedtransformerfacial] fuses the hierarchical representation extracted from CNN blocks. Its task-specific tokenization scheme facilitates multi-task learning within a single encoder architecture, effectively utilizing both self-attention and cross-attention mechanisms.
@@ 3 Methodology
Inspired by the object token in DETR [10.1007/978-3-030-58452-8_13] and mask token in MaskFormer [NEURIPS2021_950a4152, Cheng_2022_CVPR], we propose learnable knowledge tokens corresponding to learnable query features in preliminaries. DETR introduced an anchor-free semantic representation leveraging object tokens. Subsequently, Mask2Former utilized masked attention to incorporate both semantic and spatial information into learnable query features. Motivated by the concept of learnable query embeddings, we introduce a knowledge token designed to capture diverse embeddings derived from the priors of a pre-trained model. While conceptually similar to a module that fuses information acquired from an ensemble, this knowledge token is specifically tailored to the canonical task, selectively attending to and learning pertinent semantic information.
@@ 3.1 Acquiring Prior Tokens
In the preliminary stage, a set of $N$ pre-trained encoders $E=\{\varepsilon_{1},\varepsilon_{2},...,\varepsilon_{n-1},\varepsilon_{c}\}$ corresponding to the expertise of each task is prepared. These encoders could be employed either CNN or ViT architectures, or even a hybrid thereof dictated by task-specific priors. Subsequent to the preparation of $E$, a canonical task branch is selected for task-specific training. As illustrated in Fig. 3, face landmark detection serves as the canonical task in the presented example. During training stage, supervision learning is exclusively derived from the ground truth $Y=\{y_{1},y_{2},...,y_{n-1},y_{c}\}$ corresponding to the designated canonical task; ground truth of the N1N-1 remaining tasks $Y\setminus\{y_{c}\}$ is explicitly excluded from the training process.
@@
Figure 4: Detailed structure of KT-Adapter. The KT-Adapter is trained on a restricted information $T_{\text{mask}}$ to enrich the $e_{c}$ in Fig. 3. During the training phase, the MHSA layer processes $T_{\text{mask}}$, which has been subjected to subsampling via the drop mask. During inference, the MHSA layer only processes $\{t_{k},t_{c}\}$, and the canonical mask is omitted.
@@ 3.2 KT-Adapter
We introduce the KT-Adapter architecture to enable the learning of knowledge tokens. KT-Adapter functions to adapt a knowledge token, drawing inspiration from class token within attention mechanisms and object tokens in DETR [10.1007/978-3-030-58452-8_13], to a designated canonical task through the utilization of prior tokens. Leveraging a self-attention mechanism, KT-Adapter facilitates the learning of both knowledge token $t_{k}$ in KT-Adapter and canonical prior $t_{c}$ derived from canonical encoder $e_{c}$. By focusing on the canonical task, it effectively reduces redundancy among the remaining prior tokens, which represent task-optimized embeddings from pre-trained tasks, thereby generating representations $\hat{t}_{k}$ and $\hat{t}_{c}$. Distinct from ensemble and ControlNet architectures, the KT-Adapter architecture eliminates the requirement for supplementary priors during inference, thereby enabling computationally efficient inference based solely on the learned knowledge token.
@@ 3.3 Drop mask and Canonical mask
In Fig. 4, we initialize the $t_{k}$, a learnable query feature, and subsequently concatenate it with the $T$. Drawing inspiration from MaskFormer [NEURIPS2021_950a4152] and dropout [10.5555/2627435.2670313], random subsampling within the feature representation not only mitigates overfitting but also rectifies global representations and potential misalignment in the token’s forward pass between training and inference phases, specifically within KT-Adapter. We introduce a stochastic drop mask $m$ denoted as:
@@
We employed multi-head self-attention (MHSA) for learning knowledge tokens. The subsampled token set $T_{{masked}}$ in Eq. 2 is presented as input to the conventional attention mechanism:
@@ 3.4 Decoder and Inference
The $\hat{t}_{k}$ and $\hat{t}_{c}$ derived from the KT-Adapter are subjected to element-wise addition with the spatial embedding information $e_{i}$. The feature embedding $\hat{e}_{c}$ is subsequently provided as input to the canonical decoder $d_{c}$. While a set of $N$ decoders $D=\{d_{1},d_{2},...,d_{n-1},d_{c}\}$ is potentially learnable for other tasks, only the task-specific decoder $d_{c}$ undergoes fine-tuning to accommodate $\hat{e}_{c}$. This canonical, or selective, fine-tuning strategy restricts the trainable parameters to the KT-Adapter and the canonical decoder, thereby significantly reducing the computational resources required for both training and inference. Each decoder comprises a light multi-layer perceptron architecture tailored for classification and regression tasks. For the loss function, WingLoss [Feng_2018_CVPR] was employed for regression tasks, including age estimation and landmark detection, and cross-entropy loss was employed for classification tasks, including face emotion recognition and face recognition.
@@
During inference, only $t_{c}$ derived from $e_{c}$ are presented as input to the KT-Adapter as shown in Fig. 5. While the information density is lower than training, $t_{k}$ retains a sophisticated and entangled representation of the priors acquired during the training phase. The KT-Adapter’s training on ”thinned” information, facilitated by the drop mask, ensures that the inference process exhibits a high degree of fidelity to the training phase. To facilitate the application of positional embeddings to $T_{\text{masked}}$, the task index corresponding to each encoder is preserved throughout the training phase. This preservation guarantees the consistent forward pass instead of positional embeddings to corresponding token’s locations, even in scenarios where elements from $T\setminus{t_{c}}$ are absent in inference phase. Canonical masking operations are omitted during inference as it identically equals to $\{\hat{t}_{k},\hat{t}_{c}\}$. In consequence, the exclusion of other $e_{i}$ results in a substantial reduction in computational overhead compared to knowledge learning, with the computational savings scaling linearly with the number of encoders.
@@ 4 Experiments
  WFLW UTK-face Affect-Net-HQ CFP-dataset
Structure NME $\downarrow$ MAE $\downarrow$ CS@5 $\uparrow$ MAE $\downarrow$ Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$
M (baseline) 5.016 3.104 66.79 4.568 72.61 70.23 88.13 85.04
Model Ensemble (concat) 5.181 (+0.165) 3.205 (+0.100) 67.24 (+0.44%) 4.556 (-0.012) 73.59 (+0.98%) 71.16 (+0.93%) 84.93 (-3.20%) 80.91 (-4.13%)
Model Soup (Uniform) 7.864 (+2.848) 4.738 (+1.632) 50.59 (-16.2%) 6.388 (+1.820) 69.86 (-2.75%) 67.14 (-3.09%) 86.00 (-2.13%) 81.64 (-3.40%)
KT-Adapter (Ours) 4.976 (-0.040) 3.080 (-0.024) 67.29 (+0.50%) 4.548 (-0.020) 73.09 (+0.48%) 70.63 (+0.40%) 89.60 (+1.47%) 86.26 (+1.22%)
C (baseline) 4.398 2.735 68.32 4.316 74.67 72.33 93.20 91.36
Model Ensemble (concat) 4.459 (+0.061) 2.775 (+0.040) 68.60 (+0.27%) 4.285 (-0.031) 74.41 (-0.26%) 72.18 (-0.15%) 92.26 (-0.94%) 89.45 (-1.91%)
Model Soup (Uniform) 7.601 (+3.203) 4.663 (+1.928) 56.72 (-11.6%) 5.688 (+1.372) 67.63 (-7.04%) 64.56 (-7.77%) 91.73 (-1.47%) 88.97 (-2.39%)
KT-Adapter (Ours) 4.350 (-0.048) 2.708 (-0.027) 69.30 (+0.98%) 4.277 (-0.039) 74.72 (+0.05%) 72.45 (+0.12%) 94.66 (+1.46%) 92.71 (+1.35%)
S (baseline) 4.543 2.829 67.49 4.412 73.90 71.76 89.20 85.84
Model Ensemble (concat) 4.691 (+0.148) 2.912 (+0.082) 67.55 (+0.06%) 4.410 (-0.002) 75.22 (+1.32%) 72.87 (+1.11%) 89.60 (+0.40%) 85.73 (+0.11%)
Model Soup (Uniform) 7.428 (+2.885) 4.529 (+1.700) 57.58 (-9.91%) 5.477 (+1.065 70.65 (-3.25%) 67.87 (-3.89%) 89.46 (+0.26%) 85.19 (-0.65%)
KT-Adapter (Ours) 4.512 (-0.031) 2.804 (-0.025) 67.60 (+0.11%) 4.400 (-0.012) 75.20 (+1.30%) 72.77 (+1.01%) 89.73 (+0.53%) 86.44 (+0.60%)
Table 1: Comparisons to other methods. While Model Ensemble entails a time complexity of $O(N)$ both Model Soup and our proposed approach achieve a substantially lower complexity of approximately $O(1)$.
@@ 4.1 Implementation Details
The experimental evaluations were conducted on 2 NVIDIA Tesla H100 GPUs. For the backbone network architecture, MobileOne [Vasu_2023_CVPR] was selected to assess real-time performance on resource-constrained platforms, such as embedded systems and handheld devices, while ConvNeXt-b (CNN-based) [Woo_2023_CVPR] and Swin-b (ViT-based) [Liu_2021_ICCV] were employed to evaluate performance on desktop-level hardware. All input images were resized to $\times224$ pixels, with facial regions cropped using a YOLO-based [Redmon_2016_CVPR] face detector. No facial alignment procedures were applied; however, task-specific pre-processing pipelines were implemented to ensure robust generalization by tasks. Specifically, for age estimation and face recognition tasks, the facial region was cropped to encompass both the facial area and hair, whereas for face landmark detection and face emotion recognition, a more tightly cropped region was adapted. A uniform batch size of 32 was maintained across all tasks. Given the head-tuning paradigm employed, which is analogous to linear probing, a learning rate of 1e-5 was adopted in trainable layers with warm-up cosine scheduler. In terms of computational cost, the KT-Adapter exhibited 21.02M FLOPs during the inference phase, with a token dimension size of 1024. During the training phase, the computational cost increases proportionally with the number of tasks.